\begin{thebibliography}{99}
\bibitem{Boch}
S. V. Bochkarev, Everywhere divergent Fourier series with respect to the Walsh system and with respect to multiplicative systems, Russian Math. Surveys, 59(2004), no 1, 103--124.
\bibitem{GoGo}
U. Goginava and L. Gogoladze, Strong approximation by Marcinkiewicz means of two-dimensional Walsh-Fourier series, Constr. Approx., 35 (2012), no.
1, 1--19.
\bibitem{GES}
B. I. Golubov, A. V. Efimov and V. A. Skvortsov. Series and transformations of Walsh, Moscow, 1987 (Russian); English translation, Kluwer
Academic, Dordrecht, 1991.
\bibitem{FrSc1}
S. Fridli and F. Schipp, Strong summability and Sidon type inequality, Acta Sci. Math. (Szeged), 60 (1985), 277--289.
\bibitem{FrSc2}
S. Fridli and F. Schipp, Strong approximation via Sidon type inequalities, J. Approx. Theory, 94 (1998), 263--284.
\bibitem{Kar1}
G. A. Karagulyan, On the divergence of strong $\Phi$-means of Fourier series, Izv. Acad. Sci. of Armenia, 26(1991), no 2, 159--162.
\bibitem{Kar2}
G. A. Karagulyan, Everywhere divergence $\Phi$-means of Fourier series, Math. Notes, 80(2006), no 1--2, 47--56.
\bibitem{KaSa}
B. S. Kashin and A. A. Saakyan, Orthogonal series. Translated from the Russian by Ralph P. Boas. Translation edited by Ben Silver. Translations of Mathematical Monographs, 75. American Mathematical Society, Providence, RI, 1989.

\bibitem{Kol}
A. N. Kolmogoroff, Une s\'{e}rie de Fourier-Lebesque divergente presgue partout,
Comp. Rend., 183(1926), no 4,  1327--1328.
\bibitem{Kon}
S. V. Konyagin, On everywhere divergence of trigonometric Fourier series, Sb. Math., 191(2000), no 1,  97--120.
\bibitem{Mar} J.  Marcinkiewicz, Sur la sommabilit\'{e} forte des s\'{e}ries de Fourier,
J. Lond. Math. Soc., 14(1939), 162--168.
\bibitem{Osk}
K. I. Oskolkov, On strong summability of Fourier series, Trudy Mat. Inst. Steklov. 172 (1985), 280--290.
\bibitem{Rod1}
V. A. Rodin, $\BMO$-strong means of Fourier series, Functional Analysis and Its Applications
23(1989), no 2, 145--147.
\bibitem{Rod2}
V. A. Rodin, The space $\BMO$ and strong means of Walsh-Fourier series, Mathematics of the USSR-Sbornik, 74(1993), no 1, 203--218.
\bibitem{Ste}
E. M. Stein, On the limits of sequences of operators, Annals Math., 74(1961), no 2, 140--170.
\bibitem{Sch1}
F. Schipp, \"{U}ber die Divergenz der Walsh-Fourierreihen, Ann Univ. Sci. Budapest, Sec. Math., 12(1969),  49--62.
\bibitem{Sch2}
F. Schipp, \"{U}ber die Summation von Walsh-Fourierreihen,  Acta Sci. Math.(Szeged),
30(1969), 77--87.
\bibitem{SWS}
F. Schipp, W. R. Wade,  P. Simon and J. P\'{a}l, Walsh Series, an Introduction to Dyadic Harmonic Analysis, Adam Hilger, Bristol, New York, 1990.
\bibitem{Tot1} V. Totik, Notes on Fourier series strong approximations,
J. Approx. Theory, 43(1985), 105--111.
\bibitem{Tot2} V. Totik, On the strong approximation of Fourier series,
Acta Math. Acad. Sci. Hung., 1980, 1-2, 157--172.
\bibitem{Zyg} A.  Zygmund, On the convergence and summability of power series on
the circle of convergence, Proc. Lond. Math. Soc., 47(1941), 326--350.
\end{thebibliography}

\documentclass{amsart}
\usepackage{amsfonts,amsmath,amstext,amsthm,amssymb}

\newtheorem*{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\numberwithin{equation}{section}

\def\supp{{\rm supp\,}}
\def\sp{{\rm sp\,}}
\def\BMO{{\rm BMO\,}}

\def\ZI{\ensuremath{\mathbb I}}
\def\ZN{\ensuremath{\mathbb N}}
\def\ZR{\ensuremath{\mathbb R}}
\def\md#1#2\emd{\ifx0#1
\begin{equation*} #2 \end{equation*}\fi  
\ifx1#1\begin{equation}#2\end{equation}\fi   
\ifx2#1\begin{align*}#2\end{align*}\fi   
\ifx3#1\begin{align}#2\end{align}\fi    
\ifx4#1\begin{gather*}#2\end{gather*}\fi  
\ifx5#1\begin{gather}#2\end{gather}\fi   
\ifx6#1\begin{multline*}#2\end{multline*}\fi  
\ifx7#1\begin{multline}#2\end{multline}\fi  
\ifx8#1\begin{multline*}\begin{split}#2\end{split}\end{multline*}\fi
\ifx9#1\begin{multline}\begin{split}#2\end{split}\end{multline}\fi
}
\newcommand {\e }[1]{(\ref{#1})}
\newcommand {\lem }[1]{Lemma \ref{#1}}

\begin{document}
\author{G. G\'at, U. Goginava,  and G. Karagulyan}

\title{On everywhere divergence of the strong $\Phi$-means of Walsh-Fourier series}
\address{G. G\'at, Institute of Mathematics and Computer Science, College of
Ny\'\i regyh\'aza, P.O. Box 166, Nyiregyh\'aza, H-4400 Hungary}
\email{gatgy@nyf.hu}
\address{U. Goginava, Department of Mathematics, Faculty of Exact and
Natural Sciences, Tbilisi State University, Chavchavadze str. 1, Tbilisi
0128, Georgia}
\email{zazagoginava@gmail.com}

\address{G. Karagulyan, Institute of Mathematics of Armenian National Academy
of Science, Bughramian Ave. 24b, 375019, Yerevan, Armenia}
\email{g.karagulyan@yahoo.com}
\subjclass[2010]{42C10, 42A24}%
\keywords{ Walsh series, strong summability, everywhere divergent Walsh-Fourier series}
\thanks{The first author is supported by project T\'AMOP-4.2.2.A-11/1/KONV-2012-0051}

\begin{abstract}
Almost everywhere strong  exponential summability of Fourier series in Walsh and trigonometric systems established by Rodin in 1990. We prove, that if the growth of a function $\Phi(t):[0,\infty)\to[0,\infty)$ is bigger than the exponent, then the strong $\Phi$-summability of a Walsh-Fourier series can fail everywhere. The analogous theorem for trigonometric system was proved before by one of the author of this paper.
\end{abstract}
\maketitle
\section{Introduction}
In the study of almost everywhere convergence and summability  of Fourier series the trigonometric and Walsh systems have many common properties. Kolmogorov \cite{Kol} in 1926 gave a first example of everywhere divergent trigonometric Fourier series. Existence of almost everywhere divergent Walsh-Fourier series first proved by Stein \cite{Ste}. Then Schipp in \cite{Sch1} constructed an example of everywhere divergent Walsh-Fourier series. A significant complement to these divergence theorems are investigations on almost everywhere summability of Fourier series.

Let $\Phi(t):[0,\infty)\to[0,\infty)$, $\Phi(0)=0$, be an increasing continuous function. A numerical series with partial sums $s_1,s_2,\ldots $ is said to be (strong) $\Phi$-summable to a number $s$, if
\md1\label{0-1}
\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^n\Phi\left(|s_k-s|\right)=0.
\emd
We note that the condition \e{0-1} is as strong as rapidly growing is $\Phi$, and in the case of $\Phi(t)=t^p$, $p>0$, the condition \e{0-1} coincides with $H^p$-summability, well known in the theory of Fourier series. Marcinkiewicz-Zygmund in \cite{Mar}, \cite{Zyg} established almost everywhere $H^p$-summability for arbitrary trigonometric Fourier series (ordinary and conjugate). K.~I.~Oskolkov in \cite{Osk} proved a.e. $\Phi$-summability for trigonometric Fourier series if $\Phi(t)=O(t/\log\log t)$. Then V.~Rodin \cite{Rod1} established the analogous with $\Phi$ satisfying the condition
\md1\label{0-2}
\limsup_{t\to\infty}\frac{\log\Phi(t)}{t}<\infty,
\emd
which is equivalent to the bound $\Phi(t)<\exp(Ct)$ with some $C>0$. Moreover, Rodin invented an interesting  property, that is almost everywhere $\BMO$-boundedness of Fourier series, and the  a.e. $\Phi$-summability immediately follows from this results, applying John-Nirenberg theorem. G.~A.~Karagulyan in \cite{Kar1,Kar2} proved that the condition \e{0-2} is sharp for a.e. $\Phi$-summability for Fourier series. That is if
\md1\label{0-3}
\limsup_{t\to\infty}\frac{\log\Phi(t)}{t}=\infty,
\emd
then there exists an integrable function $f\in L^1(0,2\pi)$ such that
\md0
\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^n\Phi(|S_k(x,f)|)=\infty,\quad \limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^n\Phi(|\tilde S_k(x,f)|)=\infty,
\emd
holds everywhere on $\ZR$, where $S_k(x,f)$ and $\tilde S_k(x,f)$ are the ordinary and conjugate partial sums of Fourier series of $f(x)$.

Analogous problems are considered also for Walsh series. Almost everywhere $H^p$-summability of Walsh-Fourier series with $p>0$ proved by F.~Schipp in \cite{Sch2}. Almost everywhere $\Phi$-summability with condition \e{0-2} proved by V.~Rodin \cite{Rod2}.
\begin{theorem}[Rodin]If $\Phi(t):[0,\infty)\to[0,\infty)$, $\Phi(0)=0$, is an increasing continuous function satisfying \e{0-2}, then  the partial sums of Walsh-Fourier series of any function $f\in L^1[0,1)$ satisfy the condition
 \md0
\lim_{n\to\infty}\frac{1}{n}\sum_{k=1}^n\Phi\left(|S_k(x,f)-f(x)|\right)=0
\emd
almost everywhere on $[0,1)$.
\end{theorem}
In this theorem and everywhere bellow the notation $S_k(x,f)$ stands for the partial sums of Walsh-Fourier series of  $f\in L^1[0,1)$.
In the present paper we establish, that, as in trigonometric case \cite{Kar2}, the bound \e{0-2} is sharp for a.e. $\Phi$-summability of Walsh-Fourier series. Moreover, we prove
\begin{theorem}
If an increasing function $\Phi(t):[0,\infty)\to[0,\infty)$ satisfies the condition \e{0-3}, then there exists a function $f\in L^1[0,1)$ such that
\md1\label{0-4}
\limsup_{n\to\infty}\frac{1}{n}\sum_{k=1}^n\Phi\left(|S_k(x,f)|\right)=\infty
\emd
holds everywhere on $[0,1)$.
\end{theorem}
It is clear this theorem generalizes Schipp's theorem on everywhere divergence of Walsh-Fourier series.  S.~V.~Bochkarev in \cite{Boch} has constructed a function $f\in L^1[0,1)$ such that
\md1\label{0-5}
\limsup_{n\to\infty }\frac{|S_n(x,f)|}{\omega_n}=\infty
\emd
everywhere on $[0,1)$, where $\omega_n=o(\sqrt {\log n })$. It is easy to observe, that this theorem implies the existence of a function $f\in L^1[0,1)$ satisfying \e{0-4} whenever
\md0
\limsup_{t\to\infty}\frac{\log\Phi(t)}{t^2}=\infty,
\emd
instead of the condition \e{0-3}, which was the best bound before. A divergence theorem like \e{0-5} with $\omega_n=o(\sqrt {\log n/\log\log n })$ for the trigonometric Fourier series established by S.~V.~Konyagin in \cite{Kon}.

We note also, that the problem of uniformly $\Phi$-summability of trigonometric Fourier series, when $f(x)$ is a continuous function considered by V.~Totik \cite{Tot1, Tot2}. He proved that the condition \e{0-2} is necessary and sufficient for uniformly  $\Phi$-summability of Fourier series of continuous functions. For the Walsh series the analogous problem is considered by S.~Fridli and F.~Schipp \cite{FrSc1,FrSc2}, V.~Rodin \cite{Rod2},  U.~Goginava and L.~Gogoladze \cite{GoGo}.
\section{Proof of theorem}
 Recall the definitions of Rademacher and Walsh functions (see \cite{GES} or \cite{SWS}). We consider the function
\md0
r_0(x)=\left\{
\begin{array}{rcl}
1, &\hbox{ if }& x\in[0,1/2),\\
-1, &\hbox{ if }& x\in[1/2,1),
\end{array}
\right.
\emd
periodically continued over the real line. The Rademacher functions are defined by $r_k(x)=r_0(2^kx)$, $k=0,1,2,\ldots$.
Walsh system is obtained by all possible products of Rademacher functions. We shall consider the Paley ordering of Walsh system.
We set $w_0(x)\equiv 1$. To define $w_n(x)$ when $n\ge 1$ we write $n$ in dyadic form
\md1\label{a35}
n=\sum_{j=0}^k\varepsilon_j2^j,
\emd
where $\varepsilon_k=1$ and $\varepsilon_j=0$ or $1$ if $j=0,1,\ldots,k-1$, and set
\md0
w_n(x)=\prod_{j=0}^k(r_j(x))^{\varepsilon_j}.
\emd
The partial sums of Walsh-Fourier series of a function $f\in L^1[0,1)$ have a formula
\md0
S_n(x,f)=\int_0^1f(t)D_n(x\oplus t)dt,
\emd
where $D_n(x)$ is the Dirichlet kernel and $\oplus$ denotes the dyadic addition. We note that
\md0
D_{2^k}(x)=\left\{
\begin{array}{rcl}
2^k, &\hbox{ if }& x\in[0,2^{-k}),\\
0, &\hbox{ if }& x\in[2^{-k},1).
\end{array}
\right.
\emd
Dirichlet kernel can be expressed by modified Dirichlet kernel $D_n^*(x)$ by
\md0
D_n(x)=w_n(x)D_n^*(x).
\emd
If $n\in \ZN$ has the form \e{a35}, then we have
\md0
D_n^*(x)=\sum_{j=0}^k\varepsilon_jD_{2^j}^*(x)=\sum_{j=0}^k\varepsilon_jr_j(x)D_{2^j}(x).
\emd
We shall write $a\lesssim b$, if $a<c\cdot b$ and $c>0$ is an absolute constant. The notation $\ZI_E$  stands for the indicator function of a set $E$.  An interval is said to be a set of the form $[a,b)$. For a dyadic interval $\delta$ we denote by $\delta^+$ and $\delta^-$ left and right halves of $\delta$. We denote the spectrum of a Walsh polynomial $P(x)=\sum_{k=0}^m a_kw_k(x)$  by
\md0
\sp P(x)=\{k\in \ZN\cup {0}:\, a_k\neq 0\}.
\emd

In the proof of following lemma we use a well known inequality
\md1\label{a25}
\left|\left\{x\in (0,1):\, \left|\sum_{k=1}^na_kr_k(x)\right|\le \lambda\right\}\right|\ge 1-2\exp\left( -\lambda^2/4\sum_{k=1}^na_k^2\right),\quad \lambda>0,
\emd
for Rademacher polynomials (see for example \cite{KaSa}, chap. 2, theorem 5).
\begin{lemma}\label{L2}
If $n\in \ZN$, $n>50$, then there exists a set $E_n\subset [0,1)$, which is a union of some dyadic intervals of the length $2^{-n}$, satisfies
the inequality
\md1\label{a27}
|E_n|>1-2\exp(-n/36),
\emd
and for any $x\in E_n$ there exists an integer $m=m(x)<2^n$ such that 
\md1\label{a20}
\int_0^x D_m^*(x\oplus t)dt\ge \frac{n}{30}.
\emd
\end{lemma}
\begin{proof}
We define
\md1\label{a26}
E_n=\left\{x\in [0,1):\, \left|\sum_{j=1}^nr_j(x)r_{j+1}(x)\right|<\frac{n}{3}\right\}
\emd
Since $\phi_j(x)=r_j(x)r_{j+1}(x)$, $j=1,2,\ldots, n$ are independent functions, taking values $\pm 1$ equally, the inequality
\e{a25} holds for $\phi_j(x)$ functions too. Applying \e{a25} in \e{a26} we will get the bound \e{a27}. Observe that for a fixed $x\in E_n$ we have
\md1\label{a17}
\#\{j\in\ZN:\,1\le j\le n:\, r_j(x)r_{j+1}(x)=-1\}>n/3,
\emd
where $\#A$ denotes the cardinality of a set $A$. On the other hand the value in \e{a17} characterizes the number of  sign changes in the sequence $r_1(x),r_2(x),\ldots ,r_{n+1}(x)$. Using this fact, we may fix integers $1\le k_1<k_2<\ldots<k_\nu\le n$, such that
\md1\label{a18}
r_{k_i}(x)=1,\quad r_{k_i+1}(x)=-1,\quad i=1,2,\ldots , \nu,\quad \nu\ge \frac{n}{6}-1.
\emd
 Suppose $\delta_j$ is the dyadic interval of the length $2^{-j}$ containing the point $x$. Observe that \e{a18} is equivalent to the condition
\md1\label{a28}
x\in  \left(\left(\delta_{k_j}\right)^+\right)^-.
\emd
This implies
\md3
&\left(\left(\delta_{k_j}\right)^+\right)^+\subset [0,x),\label{a30}\\
&r_{k_j}(x\oplus t)= 1,\quad t\in\delta_{k_j}\cap[0,x).\label{a29}
\emd
Now consider the integer
\md0
m=2^{k_1}+2^{k_2}+\ldots+2^{k_\nu}.
\emd
Using \e{a30} and \e{a29}, we obtain
\md8
\int_0^x D_m^*(x\oplus t)dt&=\sum_{j=1}^\nu \int_0^x r_{k_j}(x\oplus t)D_{2^{k_j}}(x\oplus t)dt\\
&=\sum_{j=1}^\nu 2^{k_j}\int_{\delta_{k_j}\cap [0,x)} r_{k_j}(x\oplus t)dt\\
&\ge \sum_{j=1}^\nu 2^{k_j}\int_{\left(\left(\delta_{k_j}\right)^+\right)^+} r_{k_j}(x\oplus t)dt\\
&=\sum_{j=1}^\nu 2^{k_j-2}|\delta_{k_j}|=\frac{\nu}{4}>\frac{n}{30}.
\emd
\end{proof}
\begin{lemma}\label{L1}
For any integer $n>n_0$ there exists a Walsh polynomial $f(x)=f_n(x)$ such that
\md3
&\|f\|_1\le 4,\quad \sp f(x)\subset [p(n),q(n)],\\
\sup_{N\in [p(n),2q(n)]}&\frac{\#\{k\in \ZN:\,1\le k\le N,\, |S_k(x,f)|>n/40 \}}{N}\gtrsim 2^{-2n},\label{a32}
\emd
where $p(n)$, $q(n)$ are positive integers, and $n_0$ is an absolute constant.
\end{lemma}
\begin{proof}
We define
\md0
\theta_k=\frac{k-1}{2^n}+\frac{k-1}{4^n}\in \Delta_k=\left[\frac{k-1}{2^n},\frac{k}{2^n}\right),\quad k=1,2,\ldots ,2^n.
\emd
Let $E_n$ be the set obtained in \lem{L2}. We define $f(x)$ by
\md1\label{a11}
f(x)=2^{\gamma}\cdot \ZI_{(E_n)^c}(x)r_n(x)
+\frac{1}{2^{n}}\sum_{j=1}^{2^n}\bigg(D_{u_{2^n}}(x\oplus\theta_j)-D_{u_j}(x\oplus\theta_j)\bigg),
\emd
where
\md3
&\gamma= \big[\log_2(\exp(n/36))\big],\label{a36}\\
&u_j=2^{10 (j+n)},\quad j=1,2,\ldots , 2^n.\label{a2}
\emd
We have
\md2
&\sp\left(\ZI_{(E_n)^c}(x)r_n(x)\right)\subset [2^n,2^{n+1}),\\
&\sp\left(D_{u_{2^n}}(x\oplus\theta_j)-D_{u_j}(x\oplus\theta_j)\right)\subset (u_j,u_{2^n}]\subset [2^n,u_{2^n}],
\emd
and therefore
\md0
\sp f(x)\subset [p(n),q(n)],\quad p(n)=2^n,\quad q(n)=u_{2^n}.
\emd
Using \e{a27} and \e{a36}, we obtain
\md0
\|f\|_1\le 2^\gamma(1-|E_n|)+2\le \exp(n/36)\cdot 2\exp(-n/36)+2=  4.
\emd
From the expression \e{a11} it follows that any value taken by $f(x)$ is either $0$ or a sum of different numbers of the form
$\pm 2^k$ with $k\ge \gamma$. This implies
\md0
|f(x)|\ge 2^\gamma\ge \frac{\exp(n/36)}{2}>\frac{n}{40},\quad n>n_0=150,
\emd
whenever
\md1\label{a31}
x\in \supp f=(E_n)^c\bigcup\left(\bigcup_{j=1}^{2^n-1}(\theta_j\oplus \supp D_{u_j})\right).
\emd
On the other hand if $l\ge q(n)$ and $x$ satisfies \e{a31}, then we have
\md0
|S_l(x,f)|=|f(x)|> \frac{n}{40}.
\emd
Thus we obtain
\md0
\frac{\#\{k\in \ZN:\,1\le k\le 2q(n),\, |S_k(x,f)|>n/40 \}}{2q(n)}\ge \frac{1}{2}> 2^{-2n},
\emd
which implies \e{a32}. Now consider the case when \e{a31} doesn't hold. We may suppose that
\md1\label{a22}
x\in \Delta_k\setminus \supp f,\quad 1\le k\le 2^n.
\emd
According to \lem{L2}, there exists an integer $m=m(x)<2^n$ satisfying one of the inequality \e{a20}. First we suppose it satisfies the first one.
Together with $m$ we consider
\md0
p=p(x)=m(x)(1+2^n)<2^{2n}.
\emd
Using the definition of $\theta_j$, observe, that
\md2
&w_m(\theta_k)=w_m\left(\frac{k-1}{2^n}\right),\\
&w_{m\cdot 2^n}(\theta_k)=w_{m\cdot 2^n}\left(\frac{k-1}{4^n}\right)=w_m\left(\frac{k-1}{2^n}\right),
\emd
and therefore we get
\md1\label{a13}
w_p(\theta_k)=w_m(\theta_k)w_{m\cdot 2^n}(\theta_k)=1,\quad k=1,2,\ldots, 2^n.
\emd
Define
\md1\label{a6}
L(x)=\{l\in\ZN:\, l=p+\mu\cdot2^{2n}, \mu \in \ZN\}.
\emd
Once again using the definition of $\theta_k$ as well as \e{a13}, we conclude
\md1\label{a14}
w_l(\theta_k)=w_p(\theta_k)w_{\mu\cdot 2^{2n}}(\theta_k)=1,\quad k=1,2,\ldots,2^n,\quad l\in L(x).
\emd
Suppose
\md1\label{a16}
l\in L(x)\cap [u_{k-1},u_k),\quad k\le 2^n.
\emd
Since $x$ is taken outside of $\supp f$, we have
\md9\label{a3}
S_l(x,f)&=\frac{1}{2^{n}}\left(\sum_{j=1}^{k-1}D_l(x\oplus\theta_j)-\sum_{j=1}^{k-1}D_{u_j}(x\oplus\theta_j)\right)\\
&=\frac{1}{2^{n}}\sum_{j=1}^{k-1}D_l(x\oplus\theta_j).
\emd
On the other hand by \e{a14} we get
\md9\label{a12}
\frac{1}{2^{n}}\left|\sum_{j=1}^{k-1}D_l(x\oplus\theta_j)\right|
&=\frac{1}{2^{n}}\left|\sum_{j=1}^{k-1}w_l\left(\theta_j\right)D_l^*(x\oplus\theta_j)\right|\\
&=\frac{1}{2^{n}}\left|\sum_{j=1}^{k-1}D_l^*(x\oplus\theta_j)\right|.
\emd
Using the definition of $D_l^*(x)$, observe that
\md0
D_l^*(x)=D_p^*(x)+D_{\mu\cdot 2^{2n}}^*(x)=D_{m}^*(x)+D_{m\cdot 2^n}^*(x)+D_{\mu\cdot 2^{2n}}^*(x).
\emd
Since the supports of the functions $D_{m\cdot 2^n}^*(t)$ and $D_{\mu\cdot 2^{2n}}^*(t)$ are in $\Delta_1$, we conclude
\md1\label{a24}
D_l^*(x\oplus\theta_j)=D_m^*(x\oplus\theta_j),\quad x\in \Delta_k, \quad j\neq  k.
\emd
Thus, applying \lem{L2} and  \e{a12}, we obtain the bound
\md9\label{a15}
\frac{1}{2^{n}}\left|\sum_{j=1}^{k-1}D_l(x\oplus\theta_j)\right|=&\frac{1}{2^{n}}\left|\sum_{j=1}^{k-1}D_m^*(x\oplus\theta_j)\right|\\
\ge &\int_0^xD_m^*(x\oplus t)dt-1>\frac{n}{30}-1>\frac{n}{40},\quad n>n_0=150,
\emd
which holds whenever $l$ satisfies \e{a16}.
Taking into account of \e{a3} and \e{a15}, we get
\md0
\frac{\#\{l\in \ZN:\,1\le l\le u_k,\, |S_l(x,f)|>n/40 \}}{u_k}\ge \frac{\#\left(L(x)\cap [u_{k-1},u_k)\right)}{u_k}\gtrsim 2^{-2n},
\emd
which completes the proof of lemma.
\end{proof}
\begin{proof}[Proof of theorem]
We may choose numbers $\{n_k\}_{k=1}^\infty$ and $\{\alpha_k\}_{k=1}^\infty$ such that
\md3
&p(n_{k+1})>2q(n_k),\label{c2}\\
&\Phi\left(\frac{n_k}{50\cdot 2^k}\right)>\exp(2n_k),\label{c3}\\
&n_{k+1}>800k2^kq(n_k),\label{c4}
\emd
where $p(n)$ and $q(n)$ are the sequences determined in \lem{L1}. We just note that  \e{c3} may guarantee by using \e{0-3}. Applying \lem{L1}, we get polynomials $g_k(x)=f_{n_k}(x)$, which satisfy \e{a32} for any $x\in [0,1)$. We have
\md0
f(x)=\sum_{k=1}^\infty2^{-k}g_k(x)\in L^1[0,1).
\emd
The condition \e{c2} provides increasing spectrums of these polynomials. Thus, if $p(n_k)<l\le q(n_k)$, then we have
\md9\label{c5}
|S_l(x,f)|=&\left|\sum_{j=1}^\infty2^{-j}S_l(x,g_j)\right|=\left|\sum_{j=1}^{k-1}2^{-j}g_j(x)+2^{-k}S_l(x,g_k)\right|\\
\ge&2^{-k}|S_l(x,g_k)|-4(k-1)q(n_{k-1}).
\emd
 Applying \lem{L1}, for any $x\in[0,1)$ we may find a number $N_k\in [p(n_k),2q(n_k)]$ such  that
\md0
\#\{l\in \ZN:\,p(n_k)<l\le N_k,\, |S_l(x,g_k)|>n_k/40 \}\gtrsim \frac{N_k}{2^{2n_k}}.
\emd
Thus, using also \e{c4} and \e{c5}, we conclude
\md0
\#\{l\in \ZN:\,p(n_k)<l\le N_k,\, |S_l(x,f)|> n_k/50\cdot  2^k \}\gtrsim \frac{N_k}{2^{2n_k}}
\emd
and finally,  using \e{c3} we obtain
\md0
\frac{1}{N_k}\sum_{j=1}^{N_k}\Phi(|S_j(x,f)|)\gtrsim\frac{1}{N_k}\cdot\frac{N_k}{2^{2n_k}}\cdot\Phi\left( \frac{n_k}{50\cdot  2^k }\right)\ge \left(\frac{e}{2}\right)^{2n_k},\quad k=1,2,\ldots.
\emd
This implies the divergence of  $\Phi$-means at a point  $x\in[0,1)$ taken arbitrarily, which completes the proof of the theorem.
\end{proof}

\end{document}